A mathematics department uses 20 teaching assistants to aid in its four-semester calculus course. The number of teaching assistants assigned to each level of the course of the course depends on enrol...

  1. A mathematics department uses 20 teaching assistants to aid in its four-semester calculus course. The number of teaching assistants assigned to each level of the course of the course depends on enrollment. Here are the fall enrollments:

Calculus I

500

Calculus II

100

Calculus III

350

Calculus IV

175

Total

1125

How many teaching assistants should be assigned to teach level of course by Using the Jefferson method OR an equivalent method to Jefferson.
Answer #1

The Jefferson Method avoids the problem of an assignment resulting in a surplus or a deficit by using a divisor that will result in the correct number of assistants being assigned.

We have 1125/20 = 56.25

Calculus I : 500/56.25 = 8.88888888889

Calculus II : 100/56.25 = 1.77777777778

Calculus III : 350/56.25 = 6.22222222222

Calculus IV : 175/56.25 = 3.11111111111

Rounding them off, we get the assignment to teach level of courses as follows:

Calculus I : 9

Calculus II : 2

Calculus III : 6

Calculus IV : 3